# Diffusion approximation of the stationary distribution of a two-level single server queue

### v7

@@ 4.3 Expansion of the asymptotic BAR
${\mathcal{E}}^{(r)}_{1}+{\mathcal{E}}^{(r)}_{2}$ $=-2^{-1}\mathbb{E}^{(r)}_{s}[\{((\lambda^{(r)}_{1}\sigma^{(r)}_{e,1})^{2}+2)(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$\quad+2^{-1}\mathbb{E}^{(r)}_{s}[\{(\lambda^{(r)}_{2}(\sigma^{(r)}_{e,2})^{2}+2)(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (5.20)
We separately consider the two cases that $\lambda_{1}=\lambda_{2}$ and $\lambda_{1}\not=\lambda_{2}$. First, consider the case that $\lambda_{1}=\lambda_{2}$. In this case, $\sigma_{e,1}^{2}=\sigma_{e,2}^{2}$ by (3.f). Hence, (5.2) implies (5.2).
Next consider the case that $\lambda_{1}\not=\lambda_{2}$. In this case, we consider the BAR (4.3) for $\theta_{2}=\theta_{1}=\theta\leq 0$. From (4.2), we have
$D^{(r)}(\theta,\theta)=\alpha^{(r)}_{e}e^{r\theta(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}-e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (5.21)
Applying (4.37), (4.8) and Corollary 4.4 to this equation, we have
$D^{(r)}(\theta,\theta)$ $=\alpha^{(r)}_{e}e^{\theta\ell_{0}}(1+r\theta+o(r\theta))\mathbb{E}^{(r)}_{s}\big{[}(r\theta(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})\widehat{R}^{(r)}_{e}$
$\quad-2^{-1}r^{2}\theta^{2}\{(\lambda^{(r)}_{2})^{2}-(\lambda^{(r)}_{1})^{2}\}(\widehat{R}^{(r)}_{e})^{2})1(L^{(r)}(0)=\ell^{(r)}_{0})\big{]}+o((r\theta)^{2}),$
$=\alpha^{(r)}_{e}e^{\theta\ell_{0}}\big{\{}r\theta(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})\mathbb{E}^{(r)}_{s}\big{[}\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})\big{]}+(r\theta)^{2}({\mathcal{E}}^{(r)}_{1}+{\mathcal{E}}^{(r)}_{2})\big{\}}+o((r\theta)^{2}).$
This implies that
$D^{(r)}(\theta,\theta)$ $=r\theta\alpha^{(r)}_{e}e^{\theta\ell_{0}}(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})\mathbb{E}^{(r)}_{s}\big{[}\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})\big{]}+O((r\theta)^{2}).$ (5.22)
$\quad=c_{1}\left(\mu r\theta-\mu^{3}\sigma_{s}^{2}r^{2}\theta^{2}/2+o((\theta r)^{2})\right)\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]+D^{(r)}(\theta,\theta).$ (5.23)
Then, picking up the terms of order $r\theta$, we have
$c_{1}\mu\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]+\alpha^{(r)}_{e}e^{\theta\ell_{0}}(\lambda_{2}-\lambda_{1})\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=O(r\theta).$
$(\lambda_{2}-\lambda_{1})\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=O(r).$ (5.24)
Since $\lambda_{1}\not=\lambda_{2}$ by (f.1), (5.24) implies (5.4), which further implies (5.5). Hence,
$\lim_{r\downarrow 0}\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=0,$ (5.25)
and therefore we have (5.2) by (5.2) and Lemma 5.2. Thus, the proof of Proposition 5.2 is completed.
@@ 6.1 Process limit
Up to now, we have considered the limit of the sequence of the scaled stationary distributions in heavy traffic. From the assumptions (3.a)–(3.f), one can guess this limiting distribution must be the stationary distribution of the reflected Brownian motion on $[0,\infty)$ whose Brownian component changes its drift and variance at the level $\ell_{0}$.
@@
$\sum_{i=1}^{2}\Big{(}\!-b_{i}\theta_{i}+\frac{1}{2}c_{i}\sigma_{i}^{2}\theta_{i}^{2}\Big{)}\widetilde{\varphi}_{i}(\theta_{i})+\frac{\beta_{1}e^{\beta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}\theta_{1}\frac{c_{1}\sigma^{2}_{1}}{2}\widetilde{\varphi}_{1}(0)-\frac{\beta_{1}e^{\theta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}\theta_{1}\frac{c_{1}\sigma^{2}_{1}}{2}\widetilde{\varphi}_{1}(0)$
$\qquad+\beta_{2}e^{\theta_{2}\ell_{0}}\theta_{2}\frac{c_{2}\sigma^{2}_{2}}{2}\widetilde{\varphi}_{2}(0)=0,\qquad\theta\equiv(\theta_{1},\theta_{2})\in\mathbb{R}\times\mathbb{R}_{-},$ (6.1)
which can be obtained from (4.3) and (4.2) similarly to $D^{(r)}(\theta,\theta)$ from (5.22) and (5.2), recall that $\beta_{i}=2b_{i}/(c_{i}\sigma_{i}^{2})$ for $i=1,2$. The BAR for $b_{1}=0$ is obtained from (6.1) by letting $b_{1}\downarrow 0$. So, we only consider the case that $b_{1}\not=0$.
@@
$b(x)=b_{1}\mu 1(x\leq\ell_{0})+b_{2}\mu 1(x>\ell_{0}),$
$\sigma(x)=\sqrt{c_{1}}\mu\sigma_{1}1(x\leq\ell_{0})+\sqrt{c_{2}}\mu\sigma_{2}1(x>\ell_{0}),$
@@
$\sum_{i=1}^{2}\Big{(}\!-b_{i}\theta+\frac{1}{2}c_{i}\sigma_{i}^{2}\theta^{2}\Big{)}\widehat{\varphi}_{i}(\theta)+\frac{\beta_{1}e^{\beta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}\theta\frac{c_{1}\sigma^{2}_{1}}{2}\widehat{\varphi}_{1}(0)=0,\qquad\theta\leq 0.$ (6.3)
where $\widehat{\varphi}_{1}(\theta)=\mathbb{E}[e^{\theta Z}1(Z\leq\ell_{0})]$ and $\widehat{\varphi}_{2}(\theta)=\mathbb{E}[e^{\theta Z}1(Z>\ell_{0})]$ for random variable $Z$ subject to the stationary distribution of the $Z(\cdot)$. (6.3) is the stationary equation for $Z(\cdot)$. Then, we can see that (6.1) for $\theta_{1}=\theta_{2}=\theta$ agrees with (6.3) if and only if $c_{1}=c_{2}$, equivalently, $\lambda_{1}=\lambda_{2}$. Hence, if $L^{(r)}_{\rm dif}(\cdot)$ weakly converges to $Z(\cdot)$ as $r\downarrow 0$, then we must have $\lambda_{1}=\lambda_{2}$. Of course, this does not exclude the possibility that $L^{(r)}_{\rm dif}(\cdot)$ weakly converges to some process other than $Z(\cdot)$. We conjecture that this is indeed possible because the stationary distribution of $L^{(r)}_{\rm dif}(\cdot)$ weakly converges to the $\widetilde{\nu}$ of (3.9) for $\lambda_{1}\not=\lambda_{2}$ at least when $T_{s}$ is exponentially distribution by (ii) of Theorem 3.1. Thus, we have two challenging problems.
@@
We intuitively show (3.13). First assume that $b_{1}\not=0$. Under the stationary setting, let $\tau^{(r)}_{1}$ be the first time for $L^{(r)}(t)$ to get into $\ell^{(r)}_{0}$ from below or above. Suppose that $L^{(r)}(t)$ first moves to $\ell^{(r)}_{0}+1$ after time $\tau^{(r)}_{1}$, and returns to $\ell^{(r)}_{0}$ at time $\tau^{(r)}_{2}$ after time $B^{(r)}_{2}$. Since the relative frequency of moving up at level $\ell^{(r)}_{0}$ is guessed to be $f^{(r)}_{\rm up}\equiv\lambda^{(r)}_{2}/(\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)})$ and $B^{(r)}_{2}$ is stochastically almost identical with the busy period of the $GI/G/1$ queue with inter-arrival time $T^{(r)}_{e,2}$ and service time $T^{(r)}_{s}/c^{(r)}_{2}$, the mean sojourn time for $L^{(r)}(t)$ to be above $\ell^{(r)}_{0}$ in the cycle $(\tau^{(r)}_{1},\tau^{(r)}_{2}]$ is
@@
On the other hand, when $L^{(r)}(t)$ first moves to $\ell^{(r)}_{0}-1$ after time $\tau^{(r)}_{1}$, let $B^{(r)}_{1}$ be the time to returns to $\ell^{(r)}_{0}$ from below. Let $f^{(r)}_{\rm down}=1-f^{(r)}_{\rm up}$, then, similarly to $B^{(r)}_{2}$ but, taking $b_{1}\not=0$ and the reflecting boundary at level 00 into account, it follows from (4.25) of [12] that the mean sojourn time for $L^{(r)}(t)$ to be below $\ell^{(r)}_{0}$ in the cycle $(\tau^{(r)}_{1},\tau^{(r)}_{2}]$ is
$f^{(r)}_{\rm down}\mathbb{E}[B^{(r)}_{1}]$ $\sim\frac{c^{(r)}_{1}\mu^{(r)}}{\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)}}\times\frac{1}{\lambda^{(r)}_{1}\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{0}]}$
$=\frac{c_{1}}{c_{1}+c_{2}}\times\frac{e^{\beta_{1}\ell_{0}}-1}{rb_{1}\mu}+o(1).$ (6.5)
Since $f^{(r)}_{\rm up}B^{(r)}_{2}+f^{(r)}_{\rm down}B^{(r)}_{1}$ is the mean cycle time to return to $\ell^{(r)}_{0}$, it follows from the cycle formula of a regenerative process that
$\lim_{r\downarrow 0}\nu^{(r)}_{1}(\mathbb{R}_{+})=\lim_{r\downarrow 0}\frac{f^{(r)}_{\rm down}\mathbb{E}[B^{(r)}_{1}]}{f^{(r)}_{\rm down}\mathbb{E}[B^{(r)}_{1}]+f^{(r)}_{\rm up}\mathbb{E}[B^{(r)}_{2}]}=\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}=\widetilde{A}_{1}.$
@@
A.1 Proofs of Corollary 3.1
@@
$\frac{1}{r}\mathbb{P}[L^{(r)}=0]$ $=\frac{b_{1}e^{\beta_{1}\ell_{0}}}{c_{1}(e^{\beta_{1}\ell_{0}}-1)}\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}+o(1)$
$=\frac{b_{1}b_{2}e^{\beta_{1}\ell_{0}}}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}+o(1).$
@@
$\quad=\frac{c_{1}\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1-\beta_{1}\ell_{0})}{2b_{1}(e^{\beta_{1}\ell_{0}}-1)}\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}+\left(\ell_{0}+\frac{c_{2}\sigma_{2}^{2}}{2b_{2}}\right)\frac{c_{2}b_{1}}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}$
$\quad=\frac{c_{1}^{2}b_{2}^{2}\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)+2b_{1}b_{2}\ell_{0}(c_{2}b_{1}-c_{1}b_{2})+c_{2}^{2}b_{1}^{2}\sigma_{2}^{2}}{2b_{1}b_{2}(c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1))}.$
@@
$(\lambda_{1}+c_{1}\mu)p_{1,\ell}=\lambda_{1}p_{1,\ell-1}+c_{1}\mu p_{1,\ell+1}+\lambda_{2}p_{2,\ell-1},\quad\ell\in[1,\ell_{0}-1],$ (A.2)
$(\lambda_{1}+c_{1}\mu)p_{1,\ell_{0}}=\lambda_{1}p_{1,\ell_{0}-1}+\lambda_{2}p_{2,\ell_{0}-1},$ (A.3)
@@
$(\lambda_{2}+c_{1}\mu)p_{2,i}=c_{1}\mu p_{2,\ell+1},\qquad\ell\in[1,\ell_{0}-1],$ (A.5)
$(\lambda_{2}+c_{1}\mu)p_{2,\ell_{0}}=c_{2}\mu p_{3,\ell_{0}+1},$ (A.6)
For states in $S_{3}$, it follows from the equation for $(3,\ell_{0}+1)$ and the local balance that
$(\lambda_{2}+c_{2}\mu)p_{3,\ell_{0}+1}=\lambda_{1}p_{1,\ell_{0}}+\lambda_{2}p_{2,\ell_{0}}+c_{2}\mu p_{\ell_{0}+2},$ (A.7)
$\lambda_{2}p_{3,\ell}=c_{2}\mu p_{3,\ell+1},\qquad\ell\geq\ell_{0}+1.$ (A.8)
@@
$p_{3,\ell}=\rho_{22}^{\ell-(\ell_{0}+1)}p_{3,\ell_{0}+1},\qquad\ell\geq\ell_{0}+1.$ (A.9)
Thus, $p_{3,\ell}$ for $\ell\geq\ell_{0}+1$ is simply determined by $p_{3,\ell_{0}+1}$.
We next consider $p_{2,\ell}$ for $\ell\in[0,\ell_{0}-1]$. From (A.4) and (A.5), we have

### v8

@@ 4.3 Expansion of the asymptotic BAR
${\mathcal{E}}^{(r)}_{1}+{\mathcal{E}}^{(r)}_{2}=-2^{-1}\mathbb{E}^{(r)}_{s}[\{((\lambda^{(r)}_{1}\sigma^{(r)}_{e,1})^{2}+2)(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{1})]$
$\qquad+2^{-1}\mathbb{E}^{(r)}_{s}[\{(\lambda^{(r)}_{2}(\sigma^{(r)}_{e,2})^{2}+2)(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{1})].$ (5.19)
We separately consider the two cases that $\lambda_{1}=\lambda_{2}$ and $\lambda_{1}\not=\lambda_{2}$. First, consider the case that $\lambda_{1}=\lambda_{2}$. In this case, $\sigma_{e,1}^{2}=\sigma_{e,2}^{2}$ by (3.f). Hence, (5.2) implies (5.2). Next consider the case that $\lambda_{1}\not=\lambda_{2}$. In this case, we have (5.3) by Lemma 5.1. Applying (5.3) to (5.2), we have (5.2) because $\lambda^{(r)}_{i}$ and $\sigma^{(r)}_{e,i}$ converge to constants $\lambda_{i}$ and $\sigma_{e,i}$, respectively, as $r\downarrow 0$ for $i=1,2$. This completes the proof of Proposition 5.2.
@@ 6.1 Process limit
Up to now, we have considered the limit of the sequence of the scaled stationary distributions in heavy traffic. From the assumptions (3.a)–(3.f), one can guess this limiting distribution must be the stationary distribution of the reflected diffusion on $[0,\infty)$ whose Brownian component changes its drift and variance at the level $\ell_{1}$.
@@
$\sum_{i=1}^{2}\Big{(}\!-b_{i}\theta_{i}+\frac{1}{2}c_{i}\sigma_{i}^{2}\theta_{i}^{2}\Big{)}\widetilde{\varphi}_{i}(\theta_{i})+\frac{\beta_{1}e^{\beta_{1}\ell_{1}}}{e^{\beta_{1}\ell_{1}}-1}\theta_{1}\frac{c_{1}\sigma^{2}_{1}}{2}\widetilde{\varphi}_{1}(0)-\frac{\beta_{1}e^{\theta_{1}\ell_{1}}}{e^{\beta_{1}\ell_{1}}-1}\theta_{1}\frac{c_{1}\sigma^{2}_{1}}{2}\widetilde{\varphi}_{1}(0)$
$\qquad+\beta_{2}e^{\theta_{2}\ell_{1}}\theta_{2}\frac{c_{2}\sigma^{2}_{2}}{2}\widetilde{\varphi}_{2}(0)=0,\qquad\theta\equiv(\theta_{1},\theta_{2})\in\mathbb{R}\times\mathbb{R}_{-},$ (6.1)
which can be obtained from (4.3) and (4.2) similarly to $D^{(r)}(\theta,\theta)$ from (A.39) and (A.6), recall that $\beta_{i}=2b_{i}/(c_{i}\sigma_{i}^{2})$ for $i=1,2$. The BAR for $b_{1}=0$ is obtained from (6.1) by letting $b_{1}\downarrow 0$. So, we only consider the case that $b_{1}\not=0$.
@@
$b(x)=b_{1}\mu 1(x\leq\ell_{1})+b_{2}\mu 1(x>\ell_{1}),$
$\sigma(x)=\sqrt{c_{1}}\mu\sigma_{1}1(x\leq\ell_{1})+\sqrt{c_{2}}\mu\sigma_{2}1(x>\ell_{1}),$
@@
$\sum_{i=1}^{2}\Big{(}\!-b_{i}\theta+\frac{1}{2}c_{i}\sigma_{i}^{2}\theta^{2}\Big{)}\widehat{\varphi}_{i}(\theta)+\frac{\beta_{1}e^{\beta_{1}\ell_{1}}}{e^{\beta_{1}\ell_{1}}-1}\theta\frac{c_{1}\sigma^{2}_{1}}{2}\widehat{\varphi}_{1}(0)=0,\qquad\theta\leq 0.$ (6.3)
where $\widehat{\varphi}_{1}(\theta)=\mathbb{E}[e^{\theta Z}1(Z\leq\ell_{1})]$ and $\widehat{\varphi}_{2}(\theta)=\mathbb{E}[e^{\theta Z}1(Z>\ell_{1})]$ for random variable $Z$ subject to the stationary distribution of the $Z(\cdot)$. (6.3) is the stationary equation for $Z(\cdot)$. Then, we can see that (6.1) for $\theta_{1}=\theta_{2}=\theta$ agrees with (6.3) if and only if $c_{1}=c_{2}$, equivalently, $\lambda_{1}=\lambda_{2}$. Hence, if $L^{(r)}_{\rm dif}(\cdot)$ weakly converges to $Z(\cdot)$ as $r\downarrow 0$, then we must have $\lambda_{1}=\lambda_{2}$. Of course, this does not exclude the possibility that $L^{(r)}_{\rm dif}(\cdot)$ weakly converges to some process other than $Z(\cdot)$. We conjecture that this is indeed possible because the stationary distribution of $L^{(r)}_{\rm dif}(\cdot)$ weakly converges to the $\widetilde{\nu}$ of (3.9) for $\lambda_{1}\not=\lambda_{2}$ at least when $T_{s}$ is exponentially distribution by (ii) of Theorem 3.1. Thus, we have two challenging problems.
@@
We intuitively show (3.13). First assume that $b_{1}\not=0$. Under the stationary setting, let $\tau^{(r)}_{1}$ be the first time for $L^{(r)}(t)$ to get into $\ell^{(r)}_{1}$ from below or above. Suppose that $L^{(r)}(t)$ first moves to $\ell^{(r)}_{1}+1$ after time $\tau^{(r)}_{1}$, and returns to $\ell^{(r)}_{1}$ at time $\tau^{(r)}_{2}$ after time $B^{(r)}_{2}$. Since the relative frequency of moving up at level $\ell^{(r)}_{1}$ is guessed to be $f^{(r)}_{\rm up}\equiv\lambda^{(r)}_{2}/(\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)})$ and $B^{(r)}_{2}$ is stochastically almost identical with the busy period of the $GI/G/1$ queue with inter-arrival time $T^{(r)}_{e,2}$ and service time $T^{(r)}_{s}/c^{(r)}_{2}$, the mean sojourn time for $L^{(r)}(t)$ to be above $\ell^{(r)}_{1}$ in the cycle $(\tau^{(r)}_{1},\tau^{(r)}_{2}]$ is
@@
On the other hand, when $L^{(r)}(t)$ first moves to $\ell^{(r)}_{1}-1$ after time $\tau^{(r)}_{1}$, let $B^{(r)}_{1}$ be the time to returns to $\ell^{(r)}_{1}$ from below. Let $f^{(r)}_{\rm down}=1-f^{(r)}_{\rm up}$, then, similarly to $B^{(r)}_{2}$ but, taking $b_{1}\not=0$ and the reflecting boundary at level 00 into account, it follows from (4.25) of [12] that the mean sojourn time for $L^{(r)}(t)$ to be below $\ell^{(r)}_{1}$ in the cycle $(\tau^{(r)}_{1},\tau^{(r)}_{2}]$ is
$f^{(r)}_{\rm down}\mathbb{E}[B^{(r)}_{1}]$ $\sim\frac{c^{(r)}_{1}\mu^{(r)}}{\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)}}\times\frac{1}{\lambda^{(r)}_{1}\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{1}]}$
$=\frac{c_{1}}{c_{1}+c_{2}}\times\frac{e^{\beta_{1}\ell_{1}}-1}{rb_{1}\mu}+o(1).$ (6.5)
Since $f^{(r)}_{\rm up}B^{(r)}_{2}+f^{(r)}_{\rm down}B^{(r)}_{1}$ is the mean cycle time to return to $\ell^{(r)}_{1}$, it follows from the cycle formula of a regenerative process that
$\lim_{r\downarrow 0}\nu^{(r)}_{1}(\mathbb{R}_{+})=\lim_{r\downarrow 0}\frac{f^{(r)}_{\rm down}\mathbb{E}[B^{(r)}_{1}]}{f^{(r)}_{\rm down}\mathbb{E}[B^{(r)}_{1}]+f^{(r)}_{\rm up}\mathbb{E}[B^{(r)}_{2}]}=\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}=\widetilde{A}_{1}.$
@@
A.1 Proof of Lemma 3.1
Let $N^{(r)}_{e,\wedge}(\cdot)$ and $N^{(r)}_{e,\vee}(\cdot)$ be the renewal processes whose inter-counting times are $T_{e,1}(n)\wedge T_{e,2}(n)$ and $T_{e,1}(n)\vee T_{e,2}(n)$, respectively, for $n\geq 1$. Obviously, $0<1/\mathbb{E}[T_{e,1}\vee T_{e,2}]=\mathbb{E}[N^{(r)}_{e,\vee}(t)]\leq\mathbb{E}[N^{(r)}_{e}(t)]\leq\mathbb{E}[N^{(r)}_{e,\wedge}(t)]=1/\mathbb{E}[T_{e,1}\wedge T_{e,2}]<\infty$ under the stationary framework. Hence, $\alpha^{(r)}_{e}=\mathbb{E}[N^{(r)}_{e}(1)]$ is finite and positive. To prove $\alpha^{(r)}_{s}=\alpha^{(r)}_{e}$, consider the stationary process $X^{(r)}_{a}(t)\equiv L^{(r)}(t)\wedge a$ for each constant $a>0$. Then, $X^{(r)}_{a}(t)$ is piecewise constant, has the increment $X^{(r)}_{a}(u)-X^{(r)}_{a}(u-)=1(L^{(r)}(u-)\leq a-1)$ at the counting instant $u$ of $N^{(r)}_{e}$, and has the increment $X^{(r)}_{a}(u)-X^{(r)}_{a}(u-)=-1(L^{(r)}(u)\leq a)$ at the counting instant $u$ of $N^{(r)}_{s}$. Hence, from an elementary calculus similar to (4.1), we have
$X^{(r)}_{a}(1)-X^{(r)}_{a}(0)$ $=\int_{0}^{1}1(L^{(r)}(u-)\leq a-1)N^{(r)}_{e}(du)-\int_{0}^{1}1(L^{(r)}(u)\leq a)N^{(r)}_{s}(du).$
Since $X^{(r)}_{a}(t)$ is stationary by (3.e) and bounded by $a$, this equation yields
$\mathbb{E}\Big{[}\int_{0}^{1}1(L^{(r)}(u-)\leq a-1)N^{(r)}_{e}(du)\Big{]}=\mathbb{E}\Big{[}\int_{0}^{1}1(L^{(r)}(u)\leq a)N^{(r)}_{s}(du)\Big{]}.$
Letting $a\to\infty$, the left-hand side of this equation converges to $\mathbb{E}[N^{(r)}_{e}(1)]=\alpha^{(r)}_{e}$, while it right-hand side converges to $\mathbb{E}[N^{(r)}_{s}(1)]=\alpha^{(r)}_{s}$. Hence, we have $\alpha^{(r)}_{s}=\alpha^{(r)}_{e}$.
A.2 Proofs of Corollary 3.1
@@
$\frac{1}{r}\mathbb{P}[L^{(r)}=0]$ $=\frac{b_{1}e^{\beta_{1}\ell_{1}}}{c_{1}(e^{\beta_{1}\ell_{1}}-1)}\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}+o(1)$
$=\frac{b_{1}b_{2}e^{\beta_{1}\ell_{1}}}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}+o(1).$
@@
$\quad=\frac{c_{1}\sigma_{1}^{2}(e^{\beta_{1}\ell_{1}}-1-\beta_{1}\ell_{1})}{2b_{1}(e^{\beta_{1}\ell_{1}}-1)}\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}+\left(\ell_{1}+\frac{c_{2}\sigma_{2}^{2}}{2b_{2}}\right)\frac{c_{2}b_{1}}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}$
$\quad=\frac{c_{1}^{2}b_{2}^{2}\sigma_{1}^{2}(e^{\beta_{1}\ell_{1}}-1)+2b_{1}b_{2}\ell_{1}(c_{2}b_{1}-c_{1}b_{2})+c_{2}^{2}b_{1}^{2}\sigma_{2}^{2}}{2b_{1}b_{2}(c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1))}.$
@@
$(\lambda_{1}+c_{1}\mu)p_{1,\ell}=\lambda_{1}p_{1,\ell-1}+c_{1}\mu p_{1,\ell+1}+\lambda_{2}p_{2,\ell-1},\quad\ell\in[1,\ell_{1}-1],$ (A.2)
$(\lambda_{1}+c_{1}\mu)p_{1,\ell_{1}}=\lambda_{1}p_{1,\ell_{1}-1}+\lambda_{2}p_{2,\ell_{1}-1},$ (A.3)
@@
$(\lambda_{2}+c_{1}\mu)p_{2,i}=c_{1}\mu p_{2,\ell+1},\qquad\ell\in[1,\ell_{1}-1],$ (A.5)
$(\lambda_{2}+c_{1}\mu)p_{2,\ell_{1}}=c_{2}\mu p_{3,\ell_{1}+1},$ (A.6)
For states in $S_{3}$, it follows from the equation for $(3,\ell_{1}+1)$ and the local balance that
$(\lambda_{2}+c_{2}\mu)p_{3,\ell_{1}+1}=\lambda_{1}p_{1,\ell_{1}}+\lambda_{2}p_{2,\ell_{1}}+c_{2}\mu p_{\ell_{1}+2},$ (A.7)
$\lambda_{2}p_{3,\ell}=c_{2}\mu p_{3,\ell+1},\qquad\ell\geq\ell_{1}+1.$ (A.8)
@@
$p_{3,\ell}=\rho_{22}^{\ell-(\ell_{1}+1)}p_{3,\ell_{1}+1},\qquad\ell\geq\ell_{1}+1.$ (A.9)
Thus, $p_{3,\ell}$ for $\ell\geq\ell_{1}+1$ is simply determined by $p_{3,\ell_{1}+1}$.
We next consider $p_{2,\ell}$ for $\ell\in[0,\ell_{1}-1]$. From (A.4) and (A.5), we have